Graph the line with slope 2 passing through the point (5,3) can u please just put the points on the graph

Answer by Fombitz(32388)

Slope means change in y over change in x so a slope of 2 means you move 2 y units for every x unit.
So starting at (5,3),
(5,3)+(1,2)=(5+1,3+2)=(6,5)
(6,5) is also on the line.
.
.
.
.